### v2

# Synergy: Towards On-Body AI via Tiny AI Accelerator Collaboration on Wearables

Abstract.
The advent of tiny artificial intelligence (AI) accelerators enables AI to run at the extreme edge, offering reduced latency, lower power cost, and improved privacy. When integrated into wearable devices, these accelerators open exciting opportunities, allowing various AI apps to run directly on the body. We present Synergy that provides AI apps with best-effort performance via system-driven holistic collaboration over AI accelerator-equipped wearables. To achieve this, Synergy provides device-agnostic programming interfaces to AI apps, giving the system visibility and controllability over the app’s resource use. Then, Synergy maximizes the inference throughput of concurrent AI models by creating various execution plans for each app considering AI accelerator availability and intelligently selecting the best set of execution plans. Synergy further improves throughput by leveraging parallelization opportunities over multiple computation units. Our evaluations with 7 baselines and 8 models demonstrate that, on average, Synergy achieves a 23.0$\times$ improvement in throughput, while reducing latency by 73.9% and power consumption by 15.8%, compared to the baselines.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06copyright: none
1. Introduction
The advent of tiny artificial intelligence (AI) accelerators, such as the Analog MAX78000 (Max78000, [n. d.]), MAX78002 (Max78002, [n. d.]), Google Coral Micro (Coral Micro, [n. d.]) and GreenWaves GAP-9 (GAP8/GAP9, [n. d.]) has brought AI closer to us than ever before, offering reduced latency, low power cost, and improved privacy. These accelerators, designed for microcontrollers (MCUs) with small form factors (e.g., MAX78000: 8mm×$\times$8mm in Figure 1), are becoming integrated into wearable devices recently (Balaji and Peh, 2023; OmniBuds, [n. d.]; Shift Moonwalkers, [n. d.]; Hearables with GAP9, [n. d.]), e.g., smart earbuds, patches, watches, glasses, wristband, and shoes. This integration transforms wearable devices from smartphone-dependent accessories—used merely for data collection or notification alerts—into AI-capable devices.
With the proliferation of wearable devices, it is natural to expect that personal computing environments on the body will form a network of these AI-capable wearables. This enables a new class of on-body AI apps to emerge, which enhance app functionality by leveraging on-device AI capabilities with diverse sensing capabilities and interaction methods available across wearables. These apps, running on wearable devices, continuously monitor a variety of user contexts and proactively provide context-aware services to users directly via diverse wearable interfaces. Figure 1 shows an example scenario with three concurrent on-body AI apps: (i) memory augmentation detects greeting words using a smartwatch microphone and captures nearby faces using glasses-mounted cameras, (ii) attention alert monitors surrounding visual events through glasses and provides haptic alerts on a ring, and (iii) personal fitness coach analyzes exercise routines and vital signs on a smartwatch and smart shoes, providing auditory feedback via earbuds.
Figure 1. Wearable computing powered by tiny AI accelerators, simultaneously running on-body AI apps.
This computing environment, surrounded by AI-capable wearable devices, presents exciting opportunities for the runtime system. By leveraging the runtime collaboration of AI accelerators, it can support on-body AI apps with enhanced performance efficiently. The system can dynamically select the most suitable devices to execute tasks for sensing, model execution, and interaction. By strategically distributing model workloads across different AI accelerators, the system can prevent resource conflicts and ensure smooth operation, improving both individual app performance and overall system efficiency. Additionally, by splitting concurrent models and running the split chunks over multiple AI accelerators, the system maximizes processing capacity, supporting more models simultaneously and accommodating larger models that a single AI accelerator cannot handle.
However, it is not straightforward to realize these benefits. In the current paradigm for multi-wearable programming, task-device assignment decisions are often made at the development time. These individual app-level decisions may not be optimal at runtime due to the dynamic nature of wearable devices and dependencies on resource usage between concurrent apps. One may consider existing model partitioning techniques (Kang et al., 2017; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Jeong et al., 2018; Wang et al., 2021; Banitalebi-Dehkordi et al., 2021). While they support dynamic model splitting, they mainly focus on optimizing the splitting decision for a single AI model, thereby lacking a holistic view that considers (i) interdependency among different tasks—sensing, model inference, interaction on distributed devices—within an app and (ii) resource conflict across concurrent apps. Additionally, no existing partitioning work has been built for tiny AI accelerators yet, making it difficult to directly adopt them for on-body AI apps.
We present Synergy, a first-of-its-kind runtime system designed to efficiently support on-body AI apps on tiny AI accelerator-equipped wearable devices. At its core, Synergy provides runtime orchestration that makes holistic decisions of task-device assignments for concurrent apps over distributed AI accelerators. To this end, Synergy decouples task-device assignment decisions from app logic by providing device-agnostic programming interfaces. Then, Synergy dynamically creates various execution plans (mapping logical tasks to physical devices) for each app, considering available resources, communication over devices, model splitting options, etc. Synergy selects the optimal set of execution plans to maximize the system-wide performance. In this paper, we focus on maximizing the execution throughput of AI apps, but other objectives can be adopted as well.
To this end, Synergy has four key components. (i) Synergy generates holistic collaboration plans, each representing the mapping of all logical tasks in concurrent apps to physical resources, by holistically considering resource demands from concurrent apps and resource availability of wearables. (ii) For efficient decision-making on resource-constrained MCUs, we propose a data intensity-aware pipeline accumulation method that helps Synergy reduce the exponential search cost to linear cost while providing comparable performance to the complete search. (iii) We devise a novel latency estimation model designed for tiny AI accelerators which is used to estimate and compare the throughput of holistic collaboration plans. (vi) After deployment, Synergy further improves model throughput via an adaptive task parallelization scheduler. It reduces the latency of a selected collaboration plan by maximizing parallelization opportunities of concurrent apps at runtime across distributed computation units on wearables (e.g., processors, AI accelerators, wireless chips).
We prototyped Synergy on two tiny AI accelerator platforms, MAX78000 and MAX78002. We compare Synergy with 7 baselines including state-of-the-art model partitioning techniques (Kang et al., 2017; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Jeong et al., 2018; Wang et al., 2021; Banitalebi-Dehkordi et al., 2021). Our extensive evaluation with 8 AI models shows that Synergy consistently outperforms the baselines, with on average 23.0$\times$ throughput gain across various scenarios, while reducing latency by 73.9% and power consumption by 15.8%. Our in-depth experiments also show that Synergy effectively adapts to various runtime environment changes: the number of devices, the number of pipelines, heterogeneous device resources, different source and target mapping, and different objectives.
2. Background & Motivation
2.1. Tiny AI Accelerators
The integration of AI accelerators into MCUs represents a significant move towards distributed, on-device AI, ensuring enhanced user privacy and minimal latency. Although a number of tiny-scale accelerators have been proposed recently, very few products are commercially available and provide access and control over their underlying operations. In this paper, we chose Analog MAX78000 (Max78000, [n. d.]) and MAX78002 (Max78002, [n. d.]) as our primary platforms since they are the most widely used tiny AI accelerator research platforms (Moosmann et al., 2023; Rüegg et al., 2023; Bakar et al., 2022; Caronti et al., 2023; Moss et al., 2022) owing to the disclosed hardware details and open-source tools, enabling in-depth analysis and modification of operations.
Both MAX78000 and MAX78002 have an Arm Cortex-M4 processor with different memory capacities; 512 KB of flash and 128 KB of SRAM on MAX78000 and 2.5 MB and 384 KB on MAX78002. For acceleration, they feature a convolutional neural network (CNN) accelerator that contains 64 CNN processors specially designed for parallel convolutional operations at ultra-low power. The CNN accelerator has the dedicated memory; 512 KB of data memory, 442 KB of weight memory, and 2 KB of bias memory on MAX78000 and 1.3 MB, 2 MB, and 8 KB on MAX78002, respectively.
Figure 2. Performance comparison between AI accelerator (MAX78000) and MCUs (MAX32650 and STM32F7).
Recent benchmark study (MAX78000 Performance, [n. d.]; Moss et al., 2022) quantifies the MAX78000’s superior performance in terms of latency and energy cost. Figure 2 shows MAX78000 significantly outperforms a conventional MCU, MAX32650 with Cortex-M4 at 120 MHz (MAX32650, [n. d.]) and even a high-performance MCU, STM32F7 with Cortex-M7 at 216 MHz (STM32F7, [n. d.]) in key AI models. Latency for keyword spotting (KWS) is reduced to 2.0 ms compared to 350 ms and 123 ms for MAX32650 and STM32F7, respectively. Energy efficiency is similarly enhanced, with MAX78000 consuming merely 0.40 mJ for face detection (FaceID), in contrast to 42.1 mJ and 464 mJ consumed by MAX32650 and STM32F7.
2.2. AI Accelerator Collaboration
With wearable devices integrating tiny AI accelerators, we foresee a proliferation in on-body AI apps running across distributed wearables (Figure 1). These environments offer opportunities to enhance performance through runtime collaboration of AI accelerators as follows.
Dynamic device selection: The system can dynamically allocate tasks such as sensing, model execution, and interaction to the most suitable devices based on resource availability and conditions. For example, a higher-performance AI accelerator can be chosen for an inference task as the set of available wearable devices changes.
Strategic workload distribution: Concurrent apps often face resource conflicts. For example, assigning multiple models to the same AI accelerator can cause out-of-resource (OOR) errors if their combined size exceeds the capacity. By distributing workloads across AI accelerators, the system prevents conflicts and ensures efficient execution of apps.
Optimizing AI accelerator utilization: Tiny AI accelerators are usually optimized for single-model support. By splitting and distributing models across multiple AI accelerators, the system can maximize processing capacity. This approach supports more models simultaneously and accommodates large models that cannot fit into a single AI accelerator.
Figure 3. Comparison between (a) AI accelerator collaboration via Synergy and (b) phone offloading.
Figure 4. Comparison of Synergy and phone offloading.
2.2.1. Why not offloading?
Offloading AI model execution to a smartphone might seem advantageous due to its higher processing capability. However, this approach can be inefficient for on-body AI apps because (i) they require continuous sensing and inference for situational services and (ii) services are directly provided through wearable devices. As shown in Figure 3, offloading incurs additional data communication, resulting in increased latency and energy costs. These costs are unnecessary when apps run solely on wearables. Continuous transmission of raw sensor data to a smartphone can also cause communication and energy bottlenecks, affecting both smartphones and wearables. We compared the total throughput of model execution using Workload 1 and 2 from Section 6.2. By eliminating unnecessary links, Synergy improves throughput by 57.7$\times$ and 28.8$\times$ compared to smartphone offloading, as shown in Figure 4. Despite frequent model execution, Synergy consumes less or comparable power due to avoiding energy-intensive data communication.
3. Synergy Design
3.1. Holistic Orchestration for AI Accelerator Collaboration
3.1.1. Limitations of Existing Methods
In the current paradigm for multi-wearable programming, task-device assignments are predetermined in the app logic at development time. This app-level assignment decision has limitations in environments with concurrent AI apps on heterogeneous wearables with varying availability. For example, a memory augmentation app using a smartwatch’s AI accelerator would underperform or fail if other apps compete for the same accelerator or if the smartwatch’s battery depletes, even when other devices have available accelerators. Leveraging multiple AI accelerators is also challenging because developers lack visibility into available accelerators and their capabilities at development time.
Figure 5. Comparison of decision approaches.
Existing model partitioning techniques (Kang et al., 2017; Jeong et al., 2018; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Wang et al., 2021; Banitalebi-Dehkordi et al., 2021) designed to split AI models for offloading to powerful devices (smartphones or cloud) could be adapted for distributed on-body AI accelerators. However, these techniques primarily focus on optimizing the splitting decision for a single model. Applying these techniques to on-body AI apps presents challenges because they lack a holistic view considering (i) resource dependency among concurrent apps and (ii) interdependency between model inference and other tasks such as sensing and interaction. Additionally, no existing work is designed for partitioning models onto tiny AI accelerators, thus lacking consideration of their unique characteristics for performance modeling.
Independent decision: Existing partitioning methods split each concurrent model individually and deploy the chunks to assigned devices. This independent approach can cause resource conflicts among models, leading to performance degradation and OOR failures, especially with memory-constrained AI accelerators. For instance, as illustrated in Figure 5 (a), if two apps deploy their model chunks to the same device ($d_{1}$)—the full model of $p_{1}$’s $m_{1}$ and the partial model of $p_{2}$’s $m_{2}$—it can cause OOR on $d_{1}$ due to memory constraints.
Model-centric joint decision: Extending these approaches to jointly consider the memory capacity of split models and filter out unsupported cases can mitigate resource conflicts. However, this model-centric view still results in suboptimal performance. An app’s end-to-end pipeline includes sensing and interaction tasks before and after model inference. Data communication costs for transmitting data from the sensing device (source) to the model execution device, and then to the destination device (target), impact the app’s latency and throughput. Figure 5 (b) illustrates this. Although the split decision for $m_{1}$ and $m_{2}$ on $p_{1}$ and $p_{2}$ avoids resource conflicts, it can still cause additional communication overhead when execution is far from the source and target devices.
3.1.2. Our Approach: AI Accelerator Collaboration with a Holistic View
We propose shifting from app-level independent decision to system-driven collaboration for environments with concurrent AI apps on distributed AI accelerator-equipped wearables. The key is to decouple task-device assignment from app logic, giving the runtime system visibility and fine-grained control over AI accelerator collaboration. This allows runtime optimization of collaboration planning, leveraging direct knowledge of concurrent workloads, available resources, and dependencies. Synergy offers developers a device-agnostic programming interface (§4.2), enabling them to define the end-to-end pipeline without specifying target devices. At runtime, Synergy dynamically assigns tasks to devices to ensure best-effort performance.
The core of runtime orchestration is to exploit collaboration opportunities by (i) flexibly assigning tasks to available devices based on task requirements and device capabilities, and (ii) splitting AI models across distributed AI accelerators. We adopt a holistic view to address resource dependency between tasks within an app and between concurrent apps. For each collaboration opportunity, we abstract resource use by all concurrent apps and their tasks over distributed devices into a holistic collaboration plan. We then identify the best collaboration case based on the system objective. Figure 5 (c) shows an example of holistic collaboration by Synergy. Synergy reduces end-to-end latency by placing source and target devices closer to model execution devices on top of the best model split decision, thereby enhancing throughput.
3.2. System Scope
Moderator-initiated orchestration: In our current implementation, orchestration tasks—discovering and managing devices, creating and selecting holistic collaboration plans, and deploying them to devices—are managed by an external moderator, such as a smartphone. This is due to the limited computing capabilities of MCUs and their lower accessibility compared to smartphones, as users may not always wear their wearables. Orchestration is needed only when there is a change in apps or device availability. Once set up, runtime model inference operates independently on wearable devices. We envision a shift towards a more decentralized approach, embedding these orchestration capabilities into powerful wearables, facilitating self-sufficiency in wearable AI systems and reducing the need for external devices.
Target metric: To execute multiple models across distributed devices, various system-wide objectives can be considered, such as maximizing inference throughput, minimizing latency, and reducing energy consumption. This paper focuses on maximizing overall system-wide model inference throughput, a key quality of service metric in AI apps. To ensure fairness among models, we merge multiple app pipelines into a unified one and maximize its execution per second. Simply maximizing total throughput—the number of model executions per second—could let lower-latency models monopolize resources, causing fairness issues. This target metric can be replaced with other objectives, such as minimizing latency or energy costs. Results are reported in §6.3.4.
4. Synergy Runtime
4.1. Overview
Figure 6 illustrates the operational flow of Synergy.
Figure 6. An illustration of Synergy’s operational flow.
App development & installation stage: On-body AI apps are developed using the device-agnostic programming interfaces provided by Synergy (§4.2). Upon app installation, its end-to-end pipeline is maintained on a moderator, and tasks are dynamically distributed onto wearable devices based on orchestration decisions.
Orchestration stage: When running apps or available devices change, Synergy triggers holistic orchestration. It generates holistic collaboration plans, mapping all tasks in concurrent apps to devices (§4.3), by matching task requirements with device capabilities and splitting model tasks across distributed AI accelerators. Synergy then evaluates the potential throughput of each plan and selects the one with the maximum throughput. Two important aspects are:
  • A complete search for the optimal plan introduces intractable computation complexity, especially in resource-constrained MCUs. To reduce the search space effectively, we propose a progressive search space reduction method (§4.4), scaling down computational complexity from exponential to linear with the number of pipelines, while keeping the selected plan close to optimal.
  • Evaluating the potential throughput of candidates is crucial for plan selection. We devise a novel latency estimation model to estimate AI model execution time on tiny AI accelerators. Based on this, we build an online throughput estimation method (§4.5) to estimate the latency and throughput of a collaboration plan.
Post-deployment stage: As we combine multiple pipelines into a holistic collaboration plan, the execution cycle completes when all pipelines finish. This can lead to idle times for tasks that can be executed on parallel computation units (e.g., MCU, AI accelerator, wireless chip). To improve throughput at runtime, we propose adaptive task parallelization (§4.6), maximizing parallel execution opportunities across distributed computation units and reducing idle times.
4.2. Device-Agnostic Programming Interface
We structure on-body AI apps as a directed acyclic graph (DAG) of tasks for several reasons. First, DAGs align well with the architecture of many AI-centered apps. Second, this structure represents data and device dependencies among tasks, providing Synergy visibility and controllability of concurrent apps. Third, DAGs inherently abstract app logic as a series of tasks, simplifying the division and allocation of these tasks across distributed devices.
In this work, we categorize tasks into three types: sensing, model, and interaction. Sensing tasks are specified with requirements such as sensor type, resolution, and position. Interaction tasks are described with requirements like interface type and physical location. Currently, Synergy supports two types of requirements: designated device and sensor type for sensing tasks, and designated device or interface type for interaction tasks. For model tasks, apps specify an AI model to be executed. For example, a memory augmentation app pipeline could be (microphone, KeywordSpotting, camera on glasses), and an attention alert app could be (camera on glasses, MobileNet, haptic). In this paper, we support three tasks for a pipeline, but it can be expanded to a DAG, e.g., conditional inference or fusion models.
4.3. Holistic Collaboration Planning
Execution plan creation: As an initial step to exploit available collaboration options, Synergy creates various execution plans for each app to run over distributed devices. An execution plan abstracts the task-device assignment of a pipeline, serving as the basis for holistic orchestration. The key to obtaining diverse execution plans from each pipeline is (i) exploring various combinations of model splitting layers across distributed AI accelerators and (ii) flexibly mapping source and target tasks to suitable devices.
An execution plan is defined as a sequence of tuples, each containing a task and an assigned device (Figure 6 (b)). Synergy supports seven task types: (i) sensing, (ii) data loading to accelerator memory, (iii) (partial) model inference, (vi) data unloading from accelerator memory, (v) Tx, (vi) Rx, and (vii) interaction. For two wearables, smart glasses and a smart ring, a pipeline can be described as (camera on glasses, EfficientNet, haptic feedback); EfficientNet has 29 layers. Then, one execution plan example would be [camera $\rightarrow$ data loading $\rightarrow$ $\textit{EfficientNet}^{0:19}$rightarrow data unloading $\rightarrow$ Tx to ring] for the glasses and [Rx from glasses $\rightarrow$ data loading $\rightarrow$ $\textit{EfficientNet}^{19:29}$rightarrow data unloading $\rightarrow$ haptic] for the smart ring, where $\textit{EfficientNet}^{i:j}$ refers to the model’s layer $i$ to $j$.
Runnable holistic collaboration plan generation: The next step is to generate holistic collaboration plans, each integrating execution plans from all pipelines (one execution plan per each app). It gives Synergy visibility over resource competition and dependencies across devices (see Figure 6 (b)). A key consideration is filtering out plans that cause OOR cases. We consider three memory constraints: (i) weight memory, (ii) bias memory, and (iii) the maximum number of layers supported (e.g., 442 KB, 2 KB, 32 for MAX78000 (Max78000, [n. d.]) and 1.3 MB, 2 MB, 128 for MAX78002 (Max78002, [n. d.])). A collaboration plan is runnable if the total weight memory, bias memory, and number of layers for assigned tasks do not exceed the AI accelerators’ capacities.
4.4. Progressive Search Space Reduction
We propose a progressive search space reduction method to reduce the intractable search space for the holistic collaboration plan selection. For each pipeline $p$, the number of execution plans available $N_{p}$ can be modeled as: $N_{p}=\sum_{d=1}^{D}{}_{D}\mathrm{P}_{d}\cdot{}_{L-1}\mathrm{C}_{d-1}\cdot D^{2}$, where $D$ and $L$ denote the number of available devices and model layers, respectively. ${}_{D}\mathrm{P}_{d}$ represents possible device orders to use, ${}_{L-1}\mathrm{C}_{d-1}$ is model splitting candidates, and $D^{2}$ is all source/target device mappings, respectively. This search space grows exponentially for holistic collaboration planning with concurrent pipelines, i.e., $O(N_{p_{1}}\times N_{p_{2}}\times...)$. For instance, for three small models (a 9-layer KWS, a 14-layer SimpleNet, and a 19-layer UNet in §6.1) with three MAX78000 devices, the number of holistic collaboration plans that can be generated is $1,971\times 4,941\times 9,261=90,190,202,571$. It is evident that a complete search algorithm considering all combinations requires substantial computational overhead. To mitigate this, we propose data intensity-aware execution plan accumulation, shown in Figure 7. This approach arranges execution plans in sequence via our data intensity metric and selects an execution plan for one app at a time, building on the collaboration plan integrated with previously selected execution plans from earlier apps, thereby reducing complexity to $O(N_{p_{1}}+N_{p_{2}}+...)$.
Figure 7. Comparison between complete search (a) and our progressive pipeline selection (b).
Pipeline prioritization: The key to data intensity-aware execution plan accumulation is determining the sequence of pipelines. Synergy prioritizes data-intensive pipelines for execution plan selection, meaning pipelines that require the transmission of larger sensing data and intermediate outputs. The intuition is that optimizing data-intensive pipelines yields greater overall benefits for the system. To understand the relationship between data intensity (i.e., data size) and latency, we conducted a preliminary measurement. Figure 8 shows layer-wise latency for inference, memory, and communication, alongside output data size using UNet (Ronneberger et al., 2015) as an example. Compared to inference latency (1.5 ms), both memory latency (10.6 ms) and communication latency (6869.1 ms) are significantly higher, being 7×7$\times$7 × and 4579×4579$\times$4579 × greater, respectively. This latency gap between inference and communication is notable in tiny AI accelerators designed for fast inference. Moreover, the output size varies significantly across different layers, with the lowest layer latency at 4426.2 ms and the highest at 161864.5 ms, indicating a 36×36$\times$36 × difference. These findings suggest that (i) the end-to-end latency of a pipeline varies greatly depending on the model split in the execution plan, (ii) latency is strongly related to the size of data exchanged, and (iii) prioritizing data-intensive pipelines could offer more optimization opportunities due to fewer resource conflicts with other pipelines.
Figure 8. Layer-wise latency analysis for UNet. The y-axis is in the logarithmic scale.
Pipeline arrangement and accumulation: We define a pipeline’s data intensity as the average data size of transmission over all execution plans. Specifically, given the input size ${\tt In}^{\tt size}$ and output size ${\tt Out}^{\tt size}_{l}$ for each layer $l\in L$, the data intensity of a pipeline is defined as: $({\tt In}^{\tt size}+\sum_{l}{\tt Out}^{\tt size}_{l})/(L+1)$. The pipelines are sorted in descending order with this metric. Then, for each pipeline, Synergy generates runnable holistic collaboration plans by combining each execution plan of the current pipeline with the previously selected execution plans from earlier pipelines. Then, Synergy selects the execution plan of which holistic collaboration plan is expected to have the highest system-wide throughput based on our online throughput estimation (§4.5).
Figure 9. Comparison among different pipeline prioritization strategies against complete search (Oracle).
Comparison with other prioritization methods: We investigate (i) how significantly our selection reduces the search space compared to a complete search, and (ii) how close our prioritization method is to the optimal selection compared to other alternatives. We compare Synergy with seven baselines: the complete search (Oracle), ascending order of data intensity (DataIntensityAsc), descending/ascending order of model size (ModelSizeDes/ModelSizeAsc), and descending/ascending order of the number of model layers (NumLayersDes/NumLayersAsc), and without prioritization (Sequential). Except for Oracle, the six baselines use the execution plan accumulation method but with different prioritization objectives. In this experiment, we assume three concurrent pipelines with two MAX78000 devices and consider all possible combinations of three pipelines out of eight (Table 1). Figure 9 shows the relative throughput ratio compared to Oracle by averaging the throughput over all pipeline combinations. The results validate our design choice. Synergy outperforms the other alternatives, showing only a 3.9% degradation in overall throughput compared to the complete search. Additionally, the execution plan accumulation approach (Synergy and other prioritizations) results in a 5576$\times$ reduction in search space compared to Oracle.
4.5. Throughput Estimation for Distributed Tiny AI Accelerators
Figure 10. Tasks involved in a pipeline for latency modeling ((1)-(6)) from the viewpoint of a device.
For holistic collaboration plan selection, evaluating the throughput of a plan is necessary. To achieve this, we first model the execution latency for each task and estimate throughput based on this latency model. Figure 10 illustrates the operation of the five types of tasks ((1)-(6)) involved in a device. We present (i) our novel latency estimation model for model inference tasks, which reflects the unique characteristics of tiny AI accelerators (§4.5.1), and (ii) the latency estimation for other tasks (§4.5.2). We then introduce our technique for estimating the system-wide throughput of a holistic collaboration plan (§4.5.3).
Figure 11. Correlations of latency with (a) trainable parameters (left) and (b) accelerator clock cycles (right). Each dot means a different layer within a model.
4.5.1. Latency Modeling for Tiny AI accelerators
For latency estimation of model splitting options, existing studies rely on measurement-driven approaches: online latency measurement by running each model layer on the target device directly (Laskaridis et al., 2020; Zhang et al., 2020) or offline latency modeling with model configurations (Kang et al., 2017; Jeong et al., 2018; Wang et al., 2021), such as learning a regression model based on the number of model parameters and latency. The first approach—direct runtime measurement for each layer—is accurate and feasible for two-tier architectures (e.g., a smartphone and a cloud server) with limited cases but entails significant overhead in our scenario involving multiple models and resource-constrained devices. The second approach—latency modeling with offline profiling—mitigates this issue but has critical limitations in our environment. First, its scalability is highly limited, requiring profiling of all possible models and AI accelerators before app deployment. Second, due to the unique characteristics of AI accelerators, existing configuration-based latency modeling may not work. Figure 11 (a) shows the relationship between the number of model parameters and the corresponding inference latency of each layer in eight models on MAX78000. Due to hardware-level optimization in tiny AI accelerators, there is a weak correlation between these two variables.
To address these issues, we propose a clock cycle-based latency model. Unlike measurement-driven approaches, our model calculates the number of clock cycles required for arithmetic operations of a given layer on AI accelerators, reflecting the internal operation of convolution accelerators on tiny AI accelerators. The number of clock cycles is, by design, proportional to the latency, as the model inference task runs solely on dedicated hardware (tiny AI accelerator). This approach would be inaccurate for general-purpose processors because their use is shared with other system components or apps. Specifically, the latency within the AI accelerator can be represented as follows:
(1) $\mathcal{L}_{\tt ai\_acc}=\mathcal{L}_{\tt load}+\mathcal{L}_{\tt inf}+\mathcal{L}_{\tt unload},$
where $\mathcal{L}_{\tt inf}$ is the inference latency with the AI accelerator, and $\mathcal{L}_{\tt load}$/$\mathcal{L}_{\tt unload}$ is the data (un)loading latency between the processor (Arm Cortex-M4) and AI accelerator via SRAM.
Inference latency modeling: To model $\mathcal{L}_{\tt inf}$ ((3) in Figure 11), we calculate the number of clock cycles $\mathbb{C}$ at AI accelerators. We can estimate the latency of the inference by dividing the number of clock cycles per each layer $\mathbb{C}_{l}$ with the clock frequency $\mathcal{F}_{l}$ of the AI accelerator that processing layer $l$, i.e., $\mathcal{L}_{\tt inf}=\sum_{l}\frac{\mathbb{C}_{l}}{\mathcal{F}_{l}}$. Specifically, given the shape of an input $(H_{\tt in},W_{\tt in},C_{\tt in})$ and the corresponding output $(H_{\tt out},W_{\tt out},C_{\tt out})$ of a layer, representing the height, width, and number of channels respectively, the number of clock cycles for fully connected layers (MLP) and convolutional layers (CNN) under sequential processors such as Arm Cortex-M4 are as follows:
(2) ^$\hat{\mathbb{C}}_{\tt MLP}$ $=H_{\tt in}\cdot W_{\tt in}\cdot C_{\tt in}\cdot C_{\tt out}$
(3) ^$\hat{\mathbb{C}}_{\tt CNN}$ $=K^{2}\cdot H_{\tt in}\cdot W_{\tt out}\cdot C_{\tt in}\cdot C_{\tt out}$
$K$ refers to the size of the kernel. Note that tiny AI accelerators have (i) parallel computation units that parallelize operations across channels and (ii) convolution engines that process the convolution with a single clock cycle. Given the number of parallel convolutional processors, $P$, the clock cycles on AI accelerators are reduced to:
(4) $\mathbb{C}_{\tt MLP}$ $=H_{\tt in}\cdot W_{\tt in}\cdot\left\lceil\frac{C_{\tt in}}{P}\right\rceil\cdot C_{\tt out}$
(5) $\mathbb{C}_{\tt CNN}$ $=H_{\tt in}\cdot W_{\tt out}\cdot\left\lceil\frac{C_{\tt in}}{P}\right\rceil\cdot C_{\tt out}$
Figure 11 (b) shows the correlation between the clock cycles of every layer in eight models and the corresponding inference latency. Unlike Figure 11 (a), our clock cycle-based modeling shows a strong correlation by design. When we translate the number of clock cycles to the estimated latency with the clock frequency, the results show that the gap between the estimated and measured latency is less than 1%.
For memory operations, such as data loading and unloading between SRAM and accelerator memory ((2) and (4) in Figure 11), we use a measurement-driven approach. We estimate $\mathcal{L}_{\tt load}$ and $\mathcal{L}_{\tt unload}$ by profiling them with different data sizes and creating a linear regression model. Since data communication between two memories occurs through the central bus at a dedicated data transmission rate, the runtime latency of memory operations is proportional to the data size. As memory latency is not affected by the model architecture, we can model $\mathcal{L}_{\tt load}$ and $\mathcal{L}_{\tt unload}$ using a few data samples of different sizes.
4.5.2. Latency Modeling for Other Tasks
Besides operations on tiny AI accelerators, on-body AI apps involve sensing and communication (Rx and Tx) tasks ((1), (5), (6) in Figure 11). For sensing tasks, we measure the latency for camera and audio inputs with different parameters (e.g., different sampling rates) during the profiling phase and estimate the latency by matching the app’s requirements to the profiles. For communication latency, similar to memory operations on tiny AI accelerators, we estimate it by dividing the data size by the bandwidth of the wireless channel. Advanced methods (Kang et al., 2017; Zhang et al., 2020) can address wireless transmission latency fluctuations, but this is beyond the scope of this paper.
4.5.3. Throughput Estimation of Holistic Collaboration Plan
The next step is to estimate the throughput of a holistic collaboration plan (a collection of execution plans from multiple apps). To this end, Synergy initially estimates the end-to-end latency of a holistic collaboration plan based on the latency estimation of tasks in the plan. Since a holistic collaboration plan is structured as a DAG with multiple source tasks and target tasks, its end-to-end latency can be defined as the longest path from any source task to any target task; the path’s length is the cumulative latency of all tasks within the path. Then, the system-wide throughput is calculated by taking the inverse of the end-to-end latency and then multiplying this value by the number of pipelines.
Figure 12. Comparison between the conventional approach (a) and our proposed adaptive parallelization mechanisms (b, c) across two devices ($d_{1}$ and $d_{2}$).
4.6. Adaptive Task Parallelization
After selecting the best holistic collaboration plan, Synergy deploys the tasks and their connection to the corresponding devices. As existing model partitioning techniques (Kang et al., 2017; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Jeong et al., 2018; Wang et al., 2021; Banitalebi-Dehkordi et al., 2021) focused on a single run for a single model, a straightforward way to execute a holistic collaboration plan with concurrent pipelines over multiple runs (continuous inference over time) is to run each pipeline sequentially—immediately run another after the completion of the preceding pipeline—in a continuous manner, as shown in Figure 12 (a). However, since tasks in the plan runs over different computation units (processors, AI accelerator, and communication module) from distributed devices, such a sequential execution yields long idle time for unused units. The challenge is how to enhance performance by effectively utilizing these diverse computation units while keeping the sequential dependency of tasks within each pipeline.
To further enhance the system-wide throughput at runtime, we propose a method for adaptive task parallelization, which consists of two parallelization strategies. First, Synergy maximizes parallelization opportunities by concurrently executing tasks among different pipelines (inter-pipeline parallelization), which utilizes idle parallel computation units (Figure 12 (b)). Note that, when multiple tasks from different pipelines are competing with the same computation units, the later-arriving tasks need to wait until the completion of the earlier tasks. Second, we extend this concept to tasks from successive runs within a pipeline (inter-run parallelization) as shown in Figure 12 (c); for example, the model inference in the second run does not need to wait for the completion of the pipeline in the first run when the AI accelerator is idle and the sensor data for the second run is ready.
To enable these strategies, Synergy employs separate task queues for each type of computation unit, alongside a dedicated scheduler. At runtime, on each computation unit, its scheduler is responsible for dequeuing tasks from its queue and initiating their execution. Upon task completion, the scheduler adds the next task to its corresponding queue. This structured approach maximizes the utilization of computation units, thereby significantly improving the throughput of the holistic collaboration plan at runtime.
5. Prototype Implementation
We prototyped Synergy on the off-the-shelf MAX78000 feather board (Max78000FTHR, [n. d.]) and MAX78002 Evaluation Kit (Max78002EVKIT, [n. d.]), which are a development platform for the MAX78000 (Max78000, [n. d.]) and MAX78002 (Max78002, [n. d.]), respectively (Figure 13). Note that although the development platform is bulky, the actual size of the accelerators is tiny (e.g., 8mm×$\times$8mm for MAX78000).
Figure 13. Hardware setup.
For wireless communication, we interfaced MAX78000/ MAX78002 with an ESP8266 Wi-Fi module (ESP8266, [n. d.]). To ensure that the communication process via Wi-Fi does not disrupt the data transmissions with MAX78000/MAX78002, we have employed a round-robin scheduling algorithm on ESP8266. This mechanism is designed to alternate the data flow: transmitting data to the wireless communication channel for outward transmission and redirecting to the serial communication channel upon receiving data from the Wi-Fi interface. Additionally, we employ lightweight Internet Protocol a compact, standalone implementation of the TCP/IP protocol tailored for low-power devices, consuming a mere 40KB of memory.
The software system of Synergy is implemented on a FreeRTOS with C language. We abstract sensor reading, inference, and networking functionality as individual tasks, and each task is scheduled on top of the FreeRTOS scheduler. Once a joint plan is selected based on the target metric, the model is split to allow parts of the model to run on different AI accelerators. To this end, we synthesize the partial model for a given range of layers for a designated AI accelerator. Synthesizing is the process of generating device-specific code from pre-trained models. This involves analyzing the pre-trained model and mapping inputs, weights, and outputs to memory and processor appropriately. In this study, we generate C codes from pre-trained PyTorch models for deployment on MAX78000 and MAX78002 devices.
6. Evaluation
6.1. Experimental Setup
6.1.1. Workload
To demonstrate the effectiveness of Synergy orchestrating and executing on-body AI apps, we design four workloads with eight pipelines and eight different models (see Table 1). For the workload design, we considered two scenarios: concurrent apps, each using a different model (type, size) for different requirements (Workload 1 and 2) and a single app with a large model for enhanced capability (Workload 3 and 4). For the device setup, we used four MAX78000 devices, representing four smart wearables; an earbud, glasses, a watch, and a ring. We then assign source and target devices and a model based on the app scenarios. Figure 14 shows example execution plans of Workload 2 and 4. Pipeline 4 (keyword spotting), captures audio from an earbud ($d_{1}$), runs KWS (KWS20, [n. d.]), and sends results to a ring ($d_{4}$). Similarly, in Workload 4, pipeline 8 (object detector) captures images on glasses ($d_{2}$), runs MobileNetV2 (Sandler et al., 2018) across devices, and sends results to a ring ($d_{4}$). For each pipeline, various execution plans are generated by leveraging the four AI accelerators, including splitting models at different layers and assigning chunks to different devices.
Table 1. Workload and model description used in this study. Sizes are all in bytes.
Workload Pipeline Models Model Size Input Size Avg. Out Size
1 1 ConvNet5 71158 28$\times$28$\times$1 14031
2 ResSimpleNet(He et al., 2016) 381792 32$\times$32$\times$3 11217
3 UNet(Ronneberger et al., 2015) 279084 48$\times$48$\times$48 74547
2 4 KWS(KWS20, [n. d.]) 169472 128$\times$128$\times$1 7976
5 SimpleNet(Hasanpour et al., 2016) 166448 32$\times$32$\times$3 9237
6 WideNet(Hasanpour et al., 2016) 313700 32$\times$32$\times$3 10091
3 7 EfficientNetV2(Tan and Le, 2021) 627220 32$\times$32$\times$3 66468
4 8 MobileNetV2(Sandler et al., 2018) 821164 32$\times$32$\times$3 296318
Figure 14. Example execution plans (workload 2, 4).
Figure 15. Throughput, latency, and power consumption result for four different workloads.
6.1.2. Baselines
To the best of our knowledge, there are no existing studies that address model splitting in multi-pipeline scenarios across multiple devices. Therefore, we devised several baselines based on different rationales and adapted state-of-the-art splitting algorithms to evaluate their effectiveness in our environments. We consider 7 baselines. The first 4 are heuristic baselines that consider the resource usage of other pipelines when selecting the holistic collaboration plan. The last 3 are based on state-of-the-art algorithms.
MinDev: This heuristic aims to avoid model splitting across devices as much as possible. The rationale is that using fewer devices would reduce communication overhead between devices, thereby increasing throughput. For each pipeline, it selects an execution plan that uses the minimum number of devices to run its model, considering the resource usage of previously selected plans, similar to Synergy.
MaxDev: In contrast, MaxDev focuses on maximizing model splitting over distributed AI accelerators, with the rationale that more devices could enhance task parallelization, thereby improving throughput. Unlike MinDev, MaxDev selects an execution plan that splits the model to all available devices.
PriMinDev: Prioritized MinDev (PriMinDev) enhances MinDev by prioritizing splitting points and device assignment order. For each pipeline, it selects an execution plan that minimizes intermediate output sizes from devices, while using the fewest possible devices. When selecting the device, it prioritizes MAX78002 over MAX78000 to reduce splitting.
PriMaxDev: PriMaxDev is the same as PriMinDev except that it considers execution plans that involve all devices.
IndModel: Inspired by state-of-the-art model partitioning methods (Kang et al., 2017; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Jeong et al., 2018), IndModel selects the best-split execution plan based on metric estimations, but without a holistic view. IndModel independently selects the optimal execution plan for each pipeline and forms the final holistic collaboration plan by aggregating the selected plans.
JointModel: IndModel may lead to out-of-resource (OOR) errors if the cumulative plan exceeds available resources. To prevent this, JointModel, a multi-tenant version of IndModel, conducts a joint resource assessment similar to Synergy.
IndBest: IndBest advances IndModel by incorporating source and target devices alongside model splitting. Each pipeline independently selects the execution plan expected to yield the highest throughput by considering end-to-end latency, from sensing to output delivery. However, it does not account for the resource usage of other pipelines.
6.1.3. Performance Metrics
We consider three performance metrics: throughput, latency, and energy. Throughput is defined as the total number of model inferences per second, while latency is the time taken to execute an end-to-end holistic collaboration plan, respectively. For energy, we report the average power consumption for the execution of a holistic collaboration plan. We measured power consumption with Monsoon power monitor (Monsoon High Voltage Power Monitor, [n. d.]).
6.2. Overall Performance
Figure 15 shows the result of different methods for the four workloads. In terms of throughput, Synergy consistently outperforms all baselines owing to its holistic decision-making process and adaptive task parallelization. Synergy on average shows 23.0$\times$ higher throughput than the baselines. In Workload 1 and 2 where multiple pipelines run concurrently, IndModel (state-of-the-art model partitioning methods) results in OOR situations. In Workload 3 and 4, Synergy continues to exhibit 1.8$\times$ and 2.2$\times$ higher throughput than the second best (IndBest) respectively. The results indicate two key points. First, Synergy boosts throughput without OOR failures when multiple pipelines exist (Workload 1 and 2). Second, Synergy supports large models through the collaboration of distributed AI accelerators (Workload 3 and 4).
For the other metrics, we first found that Synergy reduces latency significantly by 73.9% on average, compared to the baselines. This is because maximizing throughput is tightly related to reduced latency; to maximize the throughput, the pipeline should be streamlined across multiple processors, which in turn reduces the end-to-end latency. Interestingly, we found that Synergy reduces power consumption by 15.8% on average, despite its significant throughput improvement of 23.0$\times$ that naturally entails more use of computation units. This is because the major source of power consumption is data transmission between devices, and maximizing throughput in Synergy naturally minimizes data communication cost.
Table 2. Ablation study of Synergy.
JRC STT PSR ATP TPUT (Wrkld 1) TPUT (Wrkld 2)
OOR OOR
0.06 2.30
0.92 15.28
2.72 15.28
4.20 29.67
6.2.1. Ablation Study
We perform an ablative study to understand the effectiveness of each technical component of Synergy. Table 2 shows the result. JRC refers to the joint resource consideration among pipelines, and STT refers to device mapping accounting for source and target tasks (§4.3). PSR means progressive search space reduction (§4.4), and ATP means adaptive task parallelization (§4.6). If none of these are applied, it is the same as the state-of-the-art approach (IndModel). If all are checked, it becomes Synergy. Overall, the throughput improves as each of our components is added. Improving IndModel by addressing resource conflicts among pipelines to avoid OOR is the same as JointModel in our baselines. Note that merely resolving the resource conflicts is still far from Synergy. Incorporating other technical components, STT, PSR, and ATP shows 6.9, 7.6, and 14.3$\times$ higher throughput compared to JRC-only, respectively.
6.3. In-Depth Analysis
6.3.1. Runtime Environment Changes
We investigate how Synergy adapts to changes in the runtime environment.
Figure 16. Impact of runtime environment changes: (a) number of devices and (b) number of pipelines.
Number of devices: In this experiment, we explore Synergy’s scalability to the change of the number of MAX78000 devices. We increase the number of devices from two to five while running the same set of four pipelines with ConvNet5, KWS, SimpleNet, and ResSimpleNet. Figure 16 (a) shows the result. Interestingly, Synergy significantly outperforms the baselines as the number of devices increases. This is due to Synergy’s strategic consideration of accelerator assignments, effectively minimizing communication overhead between source and target devices. Conversely, except for IndBest, the throughput gains for all other baselines are not notable, even with additional devices. Another interesting observation is that using more devices does not always lead to higher throughput. In the case of Synergy, the throughput saturates once the number of devices reaches 4. This is because, beyond optimal distribution, further splitting of models fails to contribute to additional throughput gains.
Number of pipelines: We assessed Synergy’s performance with a varying number of pipelines, incrementally increasing from one to six: UNet, ConvNet5, SimpleNet, KWS, ResSimpleNet, and WideNet, using four MAX78000s. To understand resource competition, we report the average throughput across pipelines, i.e., the ratio of completed pipelines per second to the total number of pipelines. Figure 16 (b) shows a downward trend in average throughput as the number of pipelines increases for all baselines due to competition for resources. Nonetheless, Synergy consistently outperforms the baselines, achieving an average throughput of 1.35 with six pipelines, 19.4$\times$ higher than the second best (PriMaxDev).
6.3.2. Composing Heterogeneous Accelerators
Figure 17. Effect of accelerator composition.
To examine Synergy’s effectiveness in heterogeneous accelerator resources, we conducted an experiment where one of four MAX78000 devices was substituted with more resource-capable MAX78002. The workload is comprised of three pipelines: ConvNet5, UNet, and EfficientNet. Figure 17 shows the results in two different setups. The inclusion of a higher-resource device generally led to an improvement in throughput. With four MAX78000s, Synergy achieves a total throughput of 0.93, outperforming the second best (JointModel) by a factor of 16.4. This throughput further increases to 3.33 when incorporating one MAX78002. PriMinDev, which allocates all models exclusively to the single MAX78002, results in a significantly low throughput of 0.06. This discrepancy highlights the importance of consideration for the communication overhead associated with source and target devices.
Figure 18. Effect of source and target mappings.
6.3.3. Source and Target Mapping
We assessed source and target mapping effects in three scenarios: (i) any device can be source or target (Any), (ii) source/target devices are evenly allocated (Distributed), and (iii) the same device is used as both source and target across pipelines (Overlapped). The Distributed setup matches Workload 1 from §6.2, while the Overlapped and Any scenarios only differ in their source and target device mappings. As shown in Figure 18, the Overlapped scenario has the lowest throughput due to communication bottlenecks, while the Any scenario has the highest throughput by distributing communication costs. The result indicates Synergy’s robustness in execution planning, considering both source and target factors.
Table 3. Comparison among different objectives.
Workload 1 TPUT (inf/s) Latency (s) Power (J/s)
TPUT-max 4.20 0.86 1.47
Latency-min 3.15 0.86 1.42
Power-min 0.19 27.17 1.22
Workload 2 TPUT (inf/s) Latency (s) Power (J/s)
TPUT-max 29.67 0.10 1.37
Latency-min 29.67 0.10 3.21
Power-min 1.37 3.21 1.06
6.3.4. Different Objectives
While Synergy defaults to maximizing system-wide throughput (TPUT-max), it can also minimize latency (Latency-min) or power consumption (Power-min). Table 3 demonstrates Synergy’s ability to achieve these different objectives effectively. Each objective excels in its metric, though identical latencies do not ensure the same holistic collaboration plan. Notably, maximizing throughput provides a balanced performance with low latency and near-minimal power use. For example, in workload 1, TPUT-max achieves 22.1$\times$ higher throughput with only 1.2$\times$ more power consumption compared to Power-min. This is because maximizing throughput reduces latency and communication overhead, the primary power consumption source.
7. Related work
Model partitioning: While there are very few attempts for partitioning AI models over multiple MCU-equipped devices, there have been active research efforts for layer-wise model partitioning over resource-limited embedded and mobile devices. In common, these methods (Kang et al., 2017; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Jeong et al., 2018; Wang et al., 2021; Banitalebi-Dehkordi et al., 2021) allocates few initial layers of a DNN on mobile/embedded device and the latter in edge or cloud server. Intermediate output from the initial layer execution is transmitted to powerful resources such as the cloud or nearby edge device where the subsequent part of the model is executed. They adapt the splitting layer depending on network status and server load. While Synergy also adopts vertical partitioning for distributed inferences, existing model partitioning studies lack the (i) end-to-end perspective of the on-body AI app execution (sensor and interface operations as well as model execution) and (ii) consideration of concurrent execution of multi-tenant models. Our experiments demonstrate that Synergy outperforms these studies by solving the challenges.
DNN workload distribution: An application pipeline often consists of a (conditional) sequence of multiple DNN models. Several distributed systems have been proposed to balance model workloads across distributed devices (Zeng et al., 2020b; Jang et al., 2021; Dong et al., 2023). Their key difference with model partitioning methods is to treat a model as a primitive execution unit and focus more on scheduling model execution over distributed devices (without model splitting). For example, a face classification task can be executed on a different device following object detection on the initial device, thereby leveraging distributed resources from multiple devices (Zeng et al., 2020b; Jang et al., 2021). Synergy shares the same high-level objective of distributing multiple model workloads into multiple devices. However, our target environment is on-body devices with AI accelerators, which brings resource challenges. To overcome the limited resources of these devices, we adopt model partitioning on top of workload distribution and devise a tailored solution for tiny AI accelerators.
Model serving systems: Several platforms, such as TensorFlow serving (TensorFlow Serving, [n. d.]), Sagemaker (SageMaker, [n. d.]), and Azure ML (ML as a Service, [n. d.]), have been proposed to facilitate model inference serving by offering containerized environments for model execution in diverse devices. Research platforms, such as Velox (Crankshaw et al., 2014) and Clipper (Crankshaw et al., 2017), focus on low-latency prediction serving, together with optimizing cloud server performance. However, in wearable computing, the main challenge lies in dynamically composing distributed tiny AI accelerators. Our device-agnostic programming interface addresses this by mapping software logic to physical resources and optimizing performance by considering pipeline interdependencies and resource usage.
8. Conclusion
We presented Synergy, a novel system that supports on-body AI apps through the collaboration of tiny AI accelerators on wearables. Synergy’s device-agnostic programming interface simplifies integrating diverse AI applications. Its runtime dynamically distributes model execution tasks to available device resources for optimal inference. Our evaluation showed Synergy achieves higher throughput than baselines.
Future work aims to enhance collaboration among AI accelerators by expanding from layer-wise model splitting to include channel-wise splitting (Mao et al., 2017a, b; Zeng et al., 2020a; Zhao et al., 2018; Hou et al., 2022; Hu and Li, 2022), enabling Synergy to handle larger inputs and optimize performance through a combination of both techniques (Pang et al., 2023). Additionally, we plan to extend Synergy to support more complex model architectures, such as sensor fusion and conditional chaining of multiple models, to meet real-world application demands. Lastly, we aim to enrich the device-agnostic programming interface by adding features like detailed taxonomy for sensor and target device mapping and supporting customized functions, thereby providing developers with greater flexibility in designing on-body AI apps.